Consider the following events in one womans life. Note that some of the events are hypothetical and some are real. Age (yeara. Calculate the total preclinical phase, in years, of this woman's cervical cancer. Assume that she did not get screened.
a. 30 years
b. 40 years
c. 10 years
d. 15 years
e. 20 years
b. Calculate the detectable preclinical phase of this woman's cervical cancer. Assume that she did not get screened.
a. 30 years
b. 10 years
c. 20 years
d. 15 years
e. 40 years
c. Calculate the lead time for this patient.
a. 20 years
b. 10 years
c. 30 years
d. 15 years
e. 60 years
Answer #1

Total preclinical phase includes biological onset of disease to symptoms of the disease.

Biological onset = 30 years

If no screening and symtomps occurs cancer detected = 60 years

60-30= 30 years of total preclinical phase

Ans a-30 year

2- detectable preclinical phase is the phase in which diagnosis should be made from the time of screening to the symptoms shows

Cancer detected by the screening = 40 years

Symtomps shows if not screen = 60 years

60-40= 20 years

Ans B- 20 years

3- lead time =The gain from screening for disease is the difference between the time a disease would have been diagnosed by clinical symptoms and when it is detected with a screening procedure. This is the lead time.

Cervical cancer begins =30 years

Cancer detect in screening = 40 years

Lead time =40-30= 10 years.

Hope you found this helpful.